LASER FREQUENCY COMBS FOR PRECISION RADIAL VELOCITY MEASUREMENTS IN ASTRO-PHYSICS

\underline{RONALD WALSWORTH}, Harvard-Smithsonian, Cambridge, MA.

Laser frequency combs adapted to serve as wavelength calibrators for astrophysical spectrographs (astro-combs) may greatly advance precision astrophysical spectroscopy, and thus provide a crucial tool for cosmology and the search for Earth-like exoplanets. I will discuss recent progress in developing astro-comb calibrators and bringing them into practice at observatories, with an emphasis on their utility for improving radial velocity precision in exoplanet studies.